simplify the expressions:
8+3(-6)/3-(-7)
1
Expert's answer
2012-06-14T07:31:11-0400
8+3(-6)/3-(-7) =8-6+7=9

though if you ment { 8+3(-6) }/{3-(-7)} (i.e. it's a fraction)
then { 8+3(-6) } / {3-(-7)}={8-18}/{3+7}=-1